OpenStudy (anonymous):

Can you please help me with some algebra? :) It's my 4.08 Polynomial Identities and Proofs assignment. Just an explanation, please!?

OpenStudy (anonymous):

Make a polynomial identity. I have to demonstrate the proof of my polynomial identity through an algebraic proof and a numerical proof in an engaging way. Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d)

OpenStudy (anonymous):

This is what i have to do using the examples provided, but i don't know how. Please help?

OpenStudy (anonymous):

Here's a polynomial identity you could make with those: (x + y)(y + x) = (x2 + 2xy + y2)

OpenStudy (anonymous):

Proof (x + y)(y + x) x(y + x) + y(y + x) distributive property xy + x2 + y2 + xy distributive x2 + xy + xy + y2 associative x2 + 2xy + y2 not sure what exactly that property is called. sorry

OpenStudy (anonymous):

For a numerical example, just plug in something for x and y and see if it works Let x = 1 and y = 2 (1 + 2) * (2 + 1) = 3 * 3 = 9 1^2 + 2*1*2 + 2^2 = 1 + 4 + 4 = 9 9 = 9 check

OpenStudy (anonymous):

Thank you so much! This was a lot of help! I will come back if i need anything else :) Thank you!!!